# SOPHY: Learning to Generate Simulation-Ready Objects with PHYsical Materials

### v2

@@ Appendix C Dataset Details
Figure 11: Statistics of fine-grained material types in our dataset.
In Table 4, we present the detailed composition of the proposed dataset, which covers $12$ common categories with part-level material annotations for $3,004$ shapes. After the VLM-guided material annotation, we obtained $8$, $7$, $10$ types of fine-grained materials for “fabric”, “leather”, and “plastic” respectively [^1] We did not obtain parts labeled with “bonded leather” for the “leather” type, and parts labeled with “flexible polyvinyl chloride”, “polystyrene” for the “plastic” type from the VLM.. The occurrence of parts labeled with these fine-grained materials is shown in Figure 11.
Furthermore, we present more dataset statistics in Figure 10. We begin by illustrating the number of rectification attempts made by experts for each object during the data verification phase outlined in Section 3.2. As presented in Figure 10(a), approximately one-third of the data has received at least one round of material parameter updates based on the experts’ feedback [^2] Please refer to Figure 13, where we provide some screenshots of the web interface we used for verifying material annotations.. This highlights the importance of human verification in correcting the initial proposals generated by the VLM. Furthermore, it is noteworthy that most verifications are completed within two iterations, which demonstrates the efficiency of our data annotation pipeline. Next, we provide the statistics of several material properties, including material behavior types in Figure 10(b), the logarithm of Young’s modulus in Figure 10(c), Poisson’s ratio in Figure 10(d), and the logarithm of yield stress in Figure 10(e). Specifically, M0 represents pure elastic materials using neo-Hookean elasticity and identity plasticity. M1 represents plastic materials with softening implemented with neo-Hookean elasticity and von Mises plasticity with damage. M2 is for plastic material without softening, which is supported by neo-Hookean elasticity and von Mises plasticity. Finally, M3 stands for granular material with StVK elasticity and Drucker-Prager plasticity as the material models. These figures showcase the material diversity of our collected dataset.
@@ Appendix D Full Prompts
In this section, we provide the detailed prompts we used in the VLM-guided material annotation stage described in Section 3.

### v3

@@ Appendix C Dataset Details
In Table˜4, we present the detailed composition of the proposed dataset, which covers $12$ common categories with part-level material annotations for $3,004$ shapes. After the VLM-guided material annotation, we obtained $8$, 77, $10$ types of fine-grained materials for “fabric”, “leather”, and “plastic” respectively [^1] We did not obtain parts labeled with “bonded leather” for the “leather” type, and parts labeled with “flexible polyvinyl chloride”, “polystyrene” for the “plastic” type from the VLM.. The occurrence of parts labeled with these fine-grained materials is shown in Figure˜11.
Furthermore, we present more dataset statistics in Figure˜12. We begin by illustrating the number of rectification attempts made by experts for each object during the data verification phase outlined in Section˜3.2. As presented in Figure˜12(a), approximately one-third of the data has received at least one round of material parameter updates based on the experts’ feedback [^2] Please refer to Figure 14, where we provide some screenshots of the web interface we used for verifying material annotations.. This highlights the importance of human verification in correcting the initial proposals generated by the VLM. Furthermore, it is noteworthy that most verifications are completed within two iterations, which demonstrates the efficiency of our data annotation pipeline. Next, we provide the statistics of several material properties, including material behavior types in Figure˜12(b), the logarithm of Young’s modulus in Figure˜12(c), Poisson’s ratio in Figure˜12(d), and the logarithm of yield stress in Figure˜12(e). Specifically, M0 represents pure elastic materials using neo-Hookean elasticity and identity plasticity. M1 represents plastic materials with softening implemented with neo-Hookean elasticity and von Mises plasticity with damage. M2 is for plastic material without softening, which is supported by neo-Hookean elasticity and von Mises plasticity. Finally, M3 stands for granular material with StVK elasticity and Drucker-Prager plasticity as the material models. These figures showcase the material diversity of our collected dataset.
@@ Appendix D Full Prompts
In this section, we provide the detailed prompts we used in the VLM-guided material annotation stage described in Section˜3.